# Stochastic SIR model with individual heterogeneity and infection-age dependent infectivity on large non-homogeneous random graphs

### v1

@@ 5.2. Convergence of $(\bar{\mu}^{I,N}_{\cdot},\bar{\mu}^{R,N}_{\cdot})$
We recall that $\mathcal{D}^{N}(t)$ is defined in (2.10) as the subset in $\mathcal{S}^{N}_{0}$ of individuals infected by the disease by time $t$. The convergence for the first term $\langle\bar{\mu}^{R,N}_{0},\varphi\rangle$ to $\langle\bar{\mu}^{R}_{0},\varphi\rangle$ follows directly from Assumption 3.1.
@@
$\mathbb{E}_{0}^{N}\Big{[}\langle\bar{\mu}^{R,N,0}_{t},\varphi\rangle\Big{]}=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\varphi(x)\bigg{(}1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg{)}\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)$ (5.45)
@@
$\langle\bar{\mu}^{R,0}_{t},\varphi\rangle=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\varphi(x)\bigg{(}1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg{)}\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)\,.$ (5.46)
@@
$\begin{split}\text{\rm Var}_{0}^{N}\Big{[}\langle\bar{\mu}^{R,N,0}_{t},\varphi\rangle\Big{]}&=\frac{1}{N}\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\varphi(x)^{2}\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\cdot\bigg{(}1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg{)}\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)\\ &\quad\leq\frac{\|\varphi\|_{\infty}^{2}}{N}\,.\end{split}$ (5.47)
@@
$\begin{split}\langle\bar{\mu}^{R,1}_{t},\varphi\rangle&=\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\varphi(x)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\cdot\exp\bigg{[}-\int_{0}^{s}\overline{\mathfrak{F}}(r,x)\mathrm{d}r\bigg{]}\bar{\mu}_{0}^{S}(\mathrm{d}x)\,\mathrm{d}s\\ &=\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\varphi(x)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\bar{\mu}_{s}^{S}(\mathrm{d}x)\,\mathrm{d}s\,.\end{split}$ (5.50)
@@
For non-negative $\varphi$, the function $:t\mapsto\langle\bar{\mu}^{R,N}_{t},\varphi\rangle$ is non-decreasing. We can thus easily adapt the argument given in Proposition 5.13, which notably involves the tightness criteria given in Proposition 5.12 with $\nu^{N}=\bar{\mu}^{R,N}$, still $\mathcal{X}=\mathbf{\mathbb{X}}$ and $\mathfrak{M}=\mathcal{C}_{b+}(\mathbf{\mathbb{X}})$ so as to apply the Second Dini theorem. For any $t$ and any compact set $K$, $\bar{\mu}^{R,N}_{t}(K^{c})\leq\bar{\mu}^{N}_{X}(K^{c})$ as defined in (3.3). Since the latter converges in probability to $\bar{\mu}_{X}$, recalling (3.3), we can find for any $\varepsilon>0$ some compact set $K_{\varepsilon}$ such that
@@
[…]
In words, $\bar{\mu}^{SI,N}$ is derived from the addition of both $\bar{\mu}^{S,N}$ and $\bar{\mu}^{I,N}$ where the first measure on $\mathbf{\mathbb{X}}$ is projected with a fixed component $-1$ according to the age variable. Intuitively, what we are doing is considering susceptible individuals as infected with infection age $-1$. The fact that $\psi$ is non-negative implies that any recovery event leads to a reduction of $\langle\bar{\mu}^{SI,N},\psi\rangle$ at this particular time. The fact that $\psi$ is non-increasing in the second variable implies that the aging of the actively infected leads as well to a reduction of $\langle\bar{\mu}^{SI,N},\psi\rangle$ over time. With this trick of combining $\bar{\mu}^{S,N}$ to $\bar{\mu}^{I,N}$ into $\bar{\mu}^{SI,N}$, any infection event leads as well to a reduction of $\langle\bar{\mu}^{SI,N},\psi\rangle$ at the infection time.
@@
The fact that the constant function equal to 1 is part of $\mathfrak{M}(\mathbf{\mathbb{X}}\times[-1,\infty))$ comes readily from the definition. If $\nu,\nu^{\prime}$ are such that $\langle\nu,\phi\rangle=\langle\nu^{\prime},\phi\rangle$ for any $\phi\in\mathfrak{M}$, then the classical approximation scheme of indicator functions by bounded continuous functions leads to the identity $\nu(A\times[-1,a])=\nu^{\prime}(A\times[-1,a])$, for any $a\in[-1,\infty)$ and measurable subset $A$ of $\mathbf{\mathbb{X}}$. The sets of this form $A\times[-1,a]$ form a $\pi$-system of subsets of the product space $\mathbf{\mathbb{X}}\times[-1,\infty)$ that contains $\mathbf{\mathbb{X}}\times[-1,\infty)$ itself and generates the Borel $\sigma$-field of $\mathbf{\mathbb{X}}\times[-1,\infty)$. The identity $\nu=\nu^{\prime}$ is thus deduced thanks e.g. to [35, Lemma 1.17]. This concludes the proof of Lemma 5.17. ∎
@@
$\mathbb{E}_{0}^{N}\Big{[}\langle\bar{\mu}^{I,N,0}_{t},\psi\rangle\Big{]}=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)$
@@
$\langle\bar{\mu}^{I,0}_{t},\psi\rangle=\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\,\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)\,.$ (5.55)
@@
$\begin{split}\text{\rm Var}_{0}^{N}\Big{[}\langle\bar{\mu}^{I,N,0}_{t},\psi\rangle\Big{]}&=\frac{1}{N}\int_{0}^{\infty}\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)^{2}\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\cdot\bigg{(}1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg{)}\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)\\ &\quad\leq\frac{\|\psi\|_{\infty}^{2}}{N}\,.\end{split}$
@@
$\begin{split}\langle\bar{\mu}^{I,1}_{t},\psi\rangle&=\int_{\mathbf{\mathbb{X}}}\int_{0}^{t}\psi(x,t-s)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\cdot\exp\bigg{[}-\int_{0}^{s}\overline{\mathfrak{F}}(r,x)\mathrm{d}r\bigg{]}\mathrm{d}s\,\bar{\mu}_{0}^{S}(\mathrm{d}x)\\ &=\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\psi(x,t-s)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\bar{\mu}_{s}^{S}(\mathrm{d}x)\,\mathrm{d}s\,.\end{split}$ (5.58)
@@ Appendix A Technical supporting results
In Appendix A, we prove three technical results that are exploited in the current paper. Lemma A.1 is used to deduce local uniform convergence in probability from pointwise estimates. Proposition A.2 is used to deduce the a.e. continuity of sections of a.e. continuous kernels. The more technical proof of proposition 5.9 is given afterwards.
@@ A.1. From pointwise to locally uniform convergence in probability
We exploit the relation between the convergence in probability and the a.s. convergence along sequence extractions as stated in [35, Lemma 4.2].
Let $(N[k])_{k\geq 1}\in\mathbb{N}^{\mathbb{N}}$ be an increasing sequence and for any $T$, let $(t_{i})_{i\geq 1}$ be a countable dense subset of $[0,T]$. By a triangular argument, we can then define an extraction $(\widetilde{N}_{n})_{n\geq 1}=(N[k_{n}])_{n\geq 1}\in\mathbb{N}^{\mathbb{N}}$, with the sequence $(k_{n})_{n\geq 1}$ being increasing, such that for any $i\geq 1$, $\psi^{\widetilde{N}_{n}}(t_{i})$ converges a.s. to $\psi(t_{i})$ as $n$ tends to infinity. Thanks to the second Dini theorem (see Exercise 127 on page 81, and its solution on page 270 in Polya and Szegö [PS1978]), this entails the a.s. convergence of $\psi^{\widetilde{N}_{n}}(t)$ to $\psi(t)$ uniformly in $t\in[0,T]$. Since the sequence $(\widetilde{N}_{n})$ is an extraction of any initial subsequence and $T$ can be freely chosen, this concludes thanks to [35, Lemma 4.2] that $\psi^{\widetilde{N}_{n}}(t)$ converges in probability to $\psi(t)$ locally uniformly in $t$. ∎
@@ A.2. Almost everywhere continuity
Proposition A.2.
Assume that $\bar{\omega}$ is $\bar{\mu}_{X}^{\otimes 2}$-a.e. continuous. Then the function $:x\mapsto\bar{\omega}(x,.)$ is $\bar{\mu}_{X}$ a.e. continuous from $\mathbf{\mathbb{X}}$ with values in $L^{1}(\bar{\mu}_{X})$.
@@
Let us define the following function of $x\in\mathbf{\mathbb{X}}$ and $\eta>0$:
$\varPsi(x;\eta)=\sup\Big{\{}\int_{\mathbf{\mathbb{X}}}|\bar{\omega}(x_{1},x^{\prime})-\bar{\omega}(x_{2},x^{\prime})|\,\bar{\mu}_{X}(\mathrm{d}x^{\prime});\,x_{1},x_{2}\in B(x;\eta)\Big{\}},$
where $B[x;\eta]$ denotes the open ball centered in $x$ of radius $\eta$. It follows from (4.6) that the irregularities of $\overline{\mathfrak{F}}$ in $x$ will be located through the following subsets $\Psi[\delta,\eta]$ of $\mathbf{\mathbb{X}}$, defined for any $\delta,\eta>0$:
$\Psi[\delta,\eta]:=\{x\in\mathbf{\mathbb{X}};\;\varPsi(x;\eta)>\delta\}.$ (A.1)
On the other hand, the $\bar{\mu}_{X}$-a.e. continuity of the kernel $\bar{\omega}$ is related to the two following similar definitions. First the function $\varPhi$ is defined for $(x,x^{\prime})\in\mathbf{\mathbb{X}}^{2}$ and $\eta>0$:
$\varPhi((x,x^{\prime});\eta)=\sup\Big{\{}|\bar{\omega}(x_{1},x^{\prime}_{1})-\bar{\omega}(x_{2},x^{\prime}_{2})|;\,(x_{1},x^{\prime}_{1}),(x_{2},x^{\prime}_{2})\in B((x,x^{\prime});\eta)\Big{\}},$
with $B[(x,x^{\prime});\eta]$ the open ball centered in $(x,x^{\prime})$ of radius $\eta$ (with the supremum norm between the two components). Then, the sets $\Phi[\delta,\eta]$ are defined for $\delta,\eta>0$ as:
$\Phi[\delta,\eta]:=\{(x,x^{\prime})\in\mathbf{\mathbb{X}}^{2};\;\varPhi((x,x^{\prime});\eta)>\delta\}.$ (A.2)
The measurability of the sets $\Phi[\delta,\eta]$ and $\Psi[\delta,\eta]$, for any $\delta,\eta>0$, is proven for completeness in Lemma A.3 just after we conclude the proof of Proposition A.2.
Note that $\cup_{\delta>0}\cap_{\eta>0}\Phi[\delta,\eta]$ is exactly the set of discontinuity points in $\mathbf{\mathbb{X}}^{2}$ of the kernel $\bar{\omega}$. $\cup_{\delta>0}\cap_{\eta>0}\Psi[\delta,\eta]$ similarly covers the set of discontinuity points in $\mathbf{\mathbb{X}}$ of the function $:x\mapsto\int\bar{\omega}(x,x^{\prime})\,\bar{\mu}_{X}(\mathrm{d}x^{\prime})$. Since $\Phi[\delta,\eta]$ is non-increasing in $\eta$ and increasing in $\delta$, the fact that $\bar{\omega}$ is $\bar{\mu}_{X}^{\otimes 2}$ a.e. continuous translates into the following convergence property to 0 for any $\delta>0$:
$\lim_{\eta\to 0}\bar{\mu}_{X}^{\otimes 2}(\Phi[\delta,\eta])=0.$ (A.3)
We will also consider the sections $\Phi_{x}[\delta,\eta]$ of $\Phi[\delta,\eta]$ with the fixed first component $x$, i.e.,
$\Phi_{x}[\delta,\eta]:=\{x^{\prime}\in\mathbf{\mathbb{X}};\,(x,x^{\prime})\in\Phi[\delta,\eta]\}.$ (A.4)
Then, by exploiting $\omega^{*}$ as the uniform upper-bound of the non-negative kernel $\bar{\omega}$ and recalling (A.2), we deduce the following inequality for any $x\in\mathbf{\mathbb{X}}$, $\delta,\eta>0$,
$\varPsi(x;\eta)\leq\omega^{*}\cdot\bar{\mu}_{X}(\Phi_{x}[\delta,\eta])+\delta.$
As a consequence, for any $x\in\Psi[2\delta,\eta]$, it holds that $\bar{\mu}_{X}(\Phi_{x}[\delta,\eta])\geq\delta/\omega^{*}$. Since $\bar{\mu}_{X}$ is a probability measure, it entails the following for any $\delta,\eta>0$,
$\bar{\mu}_{X}^{\otimes 2}(\Phi[\delta,\eta])\geq\frac{\delta}{\omega^{*}}\bar{\mu}_{X}(\Psi[2\delta,\eta]).$
Recalling from (A.3) that the l.h.s. tends to zero as $\eta$ tends to 0, so does the sequence $\bar{\mu}_{X}(\Psi[\delta,\eta])$ for any $\delta$. Up to the next Lemma A.3, this concludes the proof of Proposition A.2. ∎
@@
For any $\delta,\eta>0$, the sets $\Phi[\delta,\eta]$ and $\Psi[\delta,\eta]$ as defined respectively in (A.2) and (A.1) are measurable.

### v2

@@ 5.2. Convergence of $(\bar{\mu}^{I,N}_{\cdot},\bar{\mu}^{R,N}_{\cdot})$
We recall that $\mathcal{D}^{N}(t)$ is defined in (2.11) as the subset in $\mathcal{S}^{N}_{0}$ of individuals infected by the disease by time $t$. The convergence for the first term $\langle\bar{\mu}^{R,N}_{0},\varphi\rangle$ to $\langle\bar{\mu}^{R}_{0},\varphi\rangle$ is part of Assumption 3.1.
@@
$\mathbb{E}_{0}^{N}\Big[\langle\bar{\mu}^{R,N,0}_{t},\varphi\rangle\Big]=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\varphi(x)\bigg(1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg)\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)$ (5.45)
@@
$\langle\bar{\mu}^{R,0}_{t},\varphi\rangle=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\varphi(x)\bigg(1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg)\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)\,.$ (5.46)
@@
$\begin{split}\text{\rm Var}_{0}^{N}\Big[\langle\bar{\mu}^{R,N,0}_{t},\varphi\rangle\Big]&=\frac{1}{N}\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\varphi(x)^{2}\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\cdot\bigg(1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg)\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)\\ &\quad\leq\frac{\|\varphi\|_{\infty}^{2}}{N}\,.\end{split}$ (5.47)
@@
$\begin{split}\langle\bar{\mu}^{R,1}_{t},\varphi\rangle&=\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\varphi(x)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\cdot\exp\bigg[-\int_{0}^{s}\overline{\mathfrak{F}}(r,x)\mathrm{d}r\bigg]\bar{\mu}_{0}^{S}(\mathrm{d}x)\,\mathrm{d}s\\ &=\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\varphi(x)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\bar{\mu}_{s}^{S}(\mathrm{d}x)\,\mathrm{d}s\,.\end{split}$ (5.50)
@@
For non-negative $\varphi$, the function $t\mapsto\langle\bar{\mu}^{R,N}_{t},\varphi\rangle$ is non-decreasing. We can thus easily adapt the argument given in Proposition 5.13, which notably involves the tightness criteria given in Proposition 5.12 with $\nu^{N}=\bar{\mu}^{R,N}$, still $\mathcal{X}=\mathbf{\mathbb{X}}$ and $\mathfrak{M}=\mathcal{C}_{b+}(\mathbf{\mathbb{X}})$ so as to apply the Second Dini theorem. For any $t$ and any compact set $K$, $\bar{\mu}^{R,N}_{t}(K^{c})\leq\bar{\mu}^{N}_{X}(K^{c})$ as defined in (3.2). Since the latter converges in probability to $\bar{\mu}_{X}$, recalling (3.2), we can find for any $\varepsilon>0$ some compact set $K_{\varepsilon}$ such that
@@
$=\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\bigg(\mathbbm{1}_{\left\{D_{i}^{N}(t)=0\right\}}\psi(X_{i}^{N},-1)+\mathbbm{1}_{\left\{D_{i}^{N}(t)=1\right\}}\,\mathbbm{1}_{\left\{\eta^{N}_{i}>t-\tau_{i}^{N}\right\}}\psi(X_{i}^{N},t-\tau_{i}^{N})\bigg)$
$\quad+\frac{1}{N}\sum_{j\in\mathcal{I}^{N}(0)}\,\mathbbm{1}_{\left\{\eta_{i}^{N,0}>t\right\}}\psi(X_{j}^{N},A_{j}^{N}(0)+t)\,.$
In words, $\bar{\mu}^{SI,N}$ is derived from the addition of both $\bar{\mu}^{S,N}$ and $\bar{\mu}^{I,N}$ where the first measure on $\mathbf{\mathbb{X}}$ is projected with a fixed component 1-1 according to the age variable. Intuitively, what we are doing is considering susceptible individuals as infected with infection age 1-1. The fact that $\psi$ is non-negative implies that any recovery event leads to a reduction of $\langle\bar{\mu}^{SI,N},\psi\rangle$ at this particular time. The fact that $\psi$ is non-increasing in the second variable implies that the aging of the actively infected leads as well to a reduction of $\langle\bar{\mu}^{SI,N},\psi\rangle$ over time. With this trick of combining $\bar{\mu}^{S,N}$ to $\bar{\mu}^{I,N}$ into $\bar{\mu}^{SI,N}$, any infection event leads as well to a reduction of $\langle\bar{\mu}^{SI,N},\psi\rangle$ at the infection time.
@@
The fact that the constant function equal to 1 is part of $\mathfrak{M}(\mathbf{\mathbb{X}}\times[-1,\infty))$ comes readily from the definition. If $\nu,\nu^{\prime}$ are such that $\langle\nu,\phi\rangle=\langle\nu^{\prime},\phi\rangle$ for any $\phi\in\mathfrak{M}$, then the classical approximation scheme of indicator functions by bounded continuous functions leads to the identity $\nu(A\times[-1,a])=\nu^{\prime}(A\times[-1,a])$, for any $a\in[-1,\infty)$ and measurable subset $A$ of $\mathbf{\mathbb{X}}$. The sets of this form $A\times[-1,a]$ form a $\pi$-system of subsets of the product space $\mathbf{\mathbb{X}}\times[-1,\infty)$ that contains $\mathbf{\mathbb{X}}\times[-1,\infty)$ itself and generates the Borel $\sigma$-field of $\mathbf{\mathbb{X}}\times[-1,\infty)$. The identity $\nu=\nu^{\prime}$ is thus deduced thanks e.g. to [33, Lemma 1.17]. This concludes the proof of Lemma 5.17. ∎
@@
$\mathbb{E}_{0}^{N}\Big[\langle\bar{\mu}^{I,N,0}_{t},\psi\rangle\Big]=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)$
@@
$\langle\bar{\mu}^{I,0}_{t},\psi\rangle=\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\,\bar{\mu}_{0}^{I}(\mathrm{d}x,\mathrm{d}a)\,.$ (5.55)
@@
$\begin{split}\text{\rm Var}_{0}^{N}\Big[\langle\bar{\mu}^{I,N,0}_{t},\psi\rangle\Big]&=\frac{1}{N}\int_{0}^{\infty}\hskip-8.5359pt\int_{\mathbf{\mathbb{X}}}\psi(x,a+t)^{2}\cdot\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\cdot\bigg(1-\frac{F^{c}_{x}(a+t)}{F^{c}_{x}(a)}\bigg)\bar{\mu}_{0}^{I,N}(\mathrm{d}x,\mathrm{d}a)\\ &\quad\leq\frac{\|\psi\|_{\infty}^{2}}{N}\,.\end{split}$
@@
$\begin{split}\langle\bar{\mu}^{I,1}_{t},\psi\rangle&=\int_{\mathbf{\mathbb{X}}}\hskip-2.84544pt\int_{0}^{t}\psi(x,t-s)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\cdot\exp\bigg[-\int_{0}^{s}\overline{\mathfrak{F}}(r,x)\mathrm{d}r\bigg]\mathrm{d}s\,\bar{\mu}_{0}^{S}(\mathrm{d}x)\\ &=\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\psi(x,t-s)\cdot F_{x}(t-s)\cdot\overline{\mathfrak{F}}(s,x)\bar{\mu}_{s}^{S}(\mathrm{d}x)\,\mathrm{d}s\,.\end{split}$ (5.58)
@@ Appendix A Technical supporting results
In Appendix A, we prove three technical results that are exploited in the current paper. Lemma A.1 is used to deduce local uniform convergence in probability from pointwise estimates. Lemma A.2 is used to deduce the a.e. continuity of sections of a.e. continuous kernels and a related result is stated in the next Lemma A.3. The more technical proof of Proposition 5.9 is given afterwards.
@@ A.1. From pointwise to locally uniform convergence in probability
We exploit the relation between the convergence in probability and the a.s. convergence along sequence extractions as stated in [33, Lemma 4.2].
Let $(N[k])_{k\geq 1}\in\mathbb{N}^{\mathbb{N}}$ be an increasing sequence and for any $T$, let $(t_{i})_{i\geq 1}$ be a countable dense subset of $[0,T]$. By a triangular argument, we can then define an extraction $(\widetilde{N}_{n})_{n\geq 1}=(N[k_{n}])_{n\geq 1}\in\mathbb{N}^{\mathbb{N}}$, with the sequence $(k_{n})_{n\geq 1}$ being increasing, such that for any $i\geq 1$, $\psi^{\widetilde{N}_{n}}(t_{i})$ converges a.s. to $\psi(t_{i})$ as $n$ tends to infinity. Thanks to the second Dini theorem (see Exercise 127 on page 81, and its solution on page 270 in Polya and Szegö [46]), this entails the a.s. convergence of $\psi^{\widetilde{N}_{n}}(t)$ to $\psi(t)$ uniformly in $t\in[0,T]$. Since the sequence $(\widetilde{N}_{n})$ is an extraction of any initial subsequence and $T$ can be freely chosen, this concludes thanks to [33, Lemma 4.2] that $\psi^{\widetilde{N}_{n}}(t)$ converges in probability to $\psi(t)$ locally uniformly in $t$. ∎
@@ A.2. Almost everywhere continuity
Lemma A.2.
Let $\mathcal{X}$ and $\mathcal{Y}$ be Polish spaces. Let $\mu\in\mathcal{M}(\mathcal{X})$, $\nu\in\mathcal{M}(\mathcal{Y})$ and $k:\mathcal{X}\times\mathcal{Y}\mapsto\mathbb{R}$ be a bounded measurable function that is $(\mu\otimes\nu)$-a.e. continuous. Then the function $x\mapsto\bar{k}(x,.)$ is $\mu$ a.e. continuous from $\mathbf{\mathbb{X}}$ with values in $L^{1}(\nu)$.
Lemma A.2 is firstly exploited in the proof of Lemma 4.2 with $\mathcal{X}=\mathcal{Y}=\mathbf{\mathbb{X}}$, $\mu=\nu=\bar{\mu}_{X}$ and $k=\bar{\omega}$. It is also involved in the proof of the following Lemma A.3.
@@
For any $\delta,\eta>0$, $x\in\mathcal{X}$ and $x^{\prime}\in B(x,\eta)$:
$\int_{\mathcal{Y}}|k(x,y)-k(x^{\prime},y)|\,\nu(\mathrm{d}y)\leq\delta+\|k\|_{\infty}\,\nu\big(\mathcal{G}_{x}[\delta,\eta]),$ (A.1)
where $\mathcal{G}_{x}[\delta,\eta]=\{y\in\mathcal{Y},\,\exists x^{\prime\prime}\in B(x,\eta),\,|k(x,y)-k(x^{\prime\prime},y)|>\delta\}$.
By assumption, there exists a measurable subset $N$ of $\mathcal{X}\times\mathcal{Y}$ such that $\mu\otimes\nu(N)=0$ and such that $k$ is continuous in any $(x,y)\in\mathbf{\mathbb{X}}^{2}\setminus N$. By the Fubini-Tonelli theorem, for any $x\in\mathbf{\mathbb{X}}$, the section $N_{x}$ is measurable, where $N_{x}:=\{y\in\mathbf{\mathbb{X}},(x,y)\in N\}$. In addition, $\int_{\mathcal{X}}\nu(N_{x})\,\mu(\mathrm{d}x)=\mu\otimes\nu(N)=0$, so that $\nu(N_{x})=0$ for $\mu$ almost every $x$. For any $y\notin N_{x}$, $k$ is continuous in $(x,y)$, thus $k(x,.)$ is continuous in $y$, which entails $\cap_{n\geq 1}\mathcal{G}_{x}[\delta,2^{-n}]\in N_{x}$. For any $x$ such that $\nu(N_{x})=0$, $\nu\big(\mathcal{G}_{x}[\delta,\eta])$ tends to zero as $\eta$ tends to zero. Recalling (A.1), it concludes the proof of Lemma A.2. ∎
@@
Let $t>0$, $\mathcal{Y}$ be a Polish space and $\mu\in\mathcal{M}(\mathcal{Y})$. Let the measurable function $k:\mathcal{Y}\times[0,t]\to\mathbb{R}$ be bounded and $\mu\otimes Leb$-a.e. continuous and the measurable function $F:\mathcal{Y}\times[0,t]\to\mathbb{R}$ be bounded and such that $y\mapsto F(y,.)$ is $\mu$-a.e. continuous with respect to the uniform norm in $[0,t]$. Then, $y\mapsto\int_{0}^{t}k(y,a)\,F(y,a)\,\mathrm{d}a$ is $\mu$-a.e. continuous.
LebLeb denotes the Lebesgue measure on $[0,t]$. This lemma is to be applied in the proof of Lemma 5.10 with $\mathcal{Y}=\mathbf{\mathbb{X}}\times\mathbf{\mathbb{X}}$, $\mu(\mathrm{d}y)=\bar{\mu}_{X}\otimes\bar{\mu}_{X}(\mathrm{d}x,\mathrm{d}x^{\prime})$, $k(y,a)=\bar{\omega}(x,x^{\prime})\,\bar{\lambda}(x^{\prime},t-a)$ and $F(y,a)=\overline{\mathcal{F}}(x^{\prime},a)\cdot\exp[-\int_{0}^{a}\overline{\mathcal{F}}(x^{\prime},r)\,\mathrm{d}r]$.
It is standard that $k$ is bounded and $\mu\otimes Leb$-a.e. continuous under Assumption 2.6, 3.3 and 3.2. Thanks to (4.3), $F$ is bounded and for any $y_{1}=(x_{1},x^{\prime}_{1})\in\mathcal{Y}$ and $y_{2}=(x_{2},x^{\prime}_{2})\in\mathcal{Y}$:
$\|F(y_{1},.)-F(y_{2},.)\|_{L^{\infty}([0,t])}\leq(1+\lambda^{*}\,\omega^{*}\,t)\,\|\overline{\mathcal{F}}(x^{\prime}_{1},.)-\overline{\mathcal{F}}(x^{\prime}_{2},.)\|_{L^{\infty}([0,t])}\,$,
which entails the required regularity of $F$ thanks to Lemma 4.2.